OpenStudy (anonymous):

Simplify by expressing fractional exponents instead of radicals. ^8√(x^2y^4)

OpenStudy (solomonzelman):

? ^8√(x^2y^4)

OpenStudy (solomonzelman):

What is in front of the "^" ?

OpenStudy (anonymous):

it means it an exponent

OpenStudy (anonymous):

*its*

OpenStudy (solomonzelman):

ik, but, exponent of what?

OpenStudy (solomonzelman):

you can't just have \[^8\]

OpenStudy (anonymous):

The question that it is asking me to solve has a exponent 8√

OpenStudy (solomonzelman):

so it is just ^8 ?!

OpenStudy (solomonzelman):

maybe, \[\sqrt[8]{x^2y^4}?\]

OpenStudy (anonymous):

yes like that

OpenStudy (solomonzelman):

OK, so take them separately \[\sqrt[8]{x^2} \times \sqrt[8]{y^4}\] Now consider the following formula: \[\huge {\sqrt[a]{b^c}=b^\frac{c}{a} }\]

OpenStudy (anonymous):

im still a little confused

OpenStudy (solomonzelman):

The connection snapped, sorry, have to start over.

OpenStudy (anonymous):

its fine take your time

OpenStudy (solomonzelman):

It snapped a second time, I was almost done, I'll draw it.

OpenStudy (solomonzelman):